Can you help me please?
Graph x=3(y-1)+6
so far I have
x=3(y-1)+6
X= 3y - 3 + 6
X= 3y + 3
how do I get the y intercept so I can graph this?

Answer by Alan3354(69443)

Graph x=3(y-1)+6
so far I have
x=3(y-1)+6
X= 3y - 3 + 6
X= 3y + 3
how do I get the y intercept so I can graph this?
===========
You get the y-intercept by solving for y.
X= 3y + 3
y = (1/3)x - 1
The y-int is (0,-1) but you don't need that to graph it.
=======================
x=3(y-1)+6
Pick 2 values for y and find x
Plot the 2 points.
Draw a line thru them.
I would use y=0 and y=1